A researcher is interested in the relationship between oral contraceptive (OC) use and blood pressure in women. Suppose a sample of ten 35-to 39 year old nonpregnant, premenopausal OC users who work in a company and have a mean systolic blood pressure (SBP) of 132.86 mm Hg and sample standard deviation of 15.34 mm Hg are identified. A sample of 21 nonpregnant, premenopausal, non-OC users in the same age group are similarly identified who have mean SBP of 127. 44 mm Hg and a sample standard deviation of 18.23 mm Hg. A researcher assumes that the two groups are independent and normal will equal variances. 1. Using the p-value, does the underlying mean difference in blood pressure between the two groups differ from zero? Use the significance level α = 0.05. 2. Compute a 95% confidence interval for the true mean difference in systolic blood pressure (SBP) between 35 to 39 year old OC users and non-OC users. Interpret this result.
Answer #1

The provided sample means are shown below: X1 = 132.86 X2 = 127.44 Also, the provided sample standard deviations are: Si = 15

This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknow

t= Ã1 - X2 (ni-1)s+(n2-1) să (17+) ni+n2 - 2 0.811 132.86 – 127.44 (10-1)15.342+(21-1)18.232 (o + 11) 10+21-2 (4) Decision abFinal conclusion: Mean difference in blood pressures between the two groups not different from zero

We need to construct the 95% confidence interval for the difference between the population means Mi - M2, for the case that t

The critical value for a = 0.05 and df = 29 degrees of freedom is te = t1-a/2;n-1 = 2.045. The corresponding confidence inter

1 1 se = Sp + ni n2 1 1 = 17.385 x + 10 21 = 6.679 Now, we finally compute the confidence interval: CI = (X1 – X2 – te X se,

Final conclusion: Here we can see that within the confidence interval include zero that means, Mean difference in blood pressures between the two groups not different from zero